The areas of two similar polygons are 72 sq. cm. and 50 sq. cm. A side of the

larger polygon is 4 cm. Find the corresponding side in the smaller polygon.



1
Expert's answer
2021-02-01T08:26:50-0500

If the scale factor of the sides of two similar polygons is "k", then the ratio of the areas is "k^2" .

In our case "k^2=\frac{72}{50}=1.44."

So, "k=\sqrt{1.44}=1.2."

The corresponding side in the smaller polygon: "\frac{4}{1.2}=\frac{10}{3}" cm.